OpenStudy (anonymous):

calculate the value of y for which 2log(3)y-log(3)(y-4)=2 (the 3 is the base in both cases i just didnt know how to show subscript on my computer)

OpenStudy (freckles):

\[2 \log_3(y)-\log_3(y-4)=2\]

OpenStudy (freckles):

use power rule first for the 2log(y) thing

OpenStudy (freckles):

then you will want to use quotient rule for the the part that results from that

OpenStudy (freckles):

useful rules here: \[r \log_a(y)=\log_a(y^r) \text{ power rule } \\ \log_a(x)-\log_a(y)=\log_a(\frac{x}{y}) \text{ quotient rule } \\ \log_a(w)=q \text{ implies } a^q=w \]

OpenStudy (anonymous):

ok so i have log(3)(y^2/y+4)=2

OpenStudy (anonymous):

whats next?

OpenStudy (freckles):

the last thing I wrote is next

OpenStudy (freckles):

\[\log_a(w)=q \text{ implies } a^q=w\]

OpenStudy (freckles):

and i think you meant \[\log_3(\frac{y^2}{y-4})=2\]

OpenStudy (freckles):

but anyways notice that the input becomes the output and the output becomes the input when using the equivalent form I just mentioned

OpenStudy (freckles):

that is \[\log_3(\frac{y^2}{y-4})=2 \text{ implies } 3^{2}=\frac{y^2}{y-4} \\ \text{ just used } \log_a(w)=q \text{ implies } a^q=w\]

OpenStudy (anonymous):

oh great thanks i got it to a quadratic of y^2-9y-36=0 (i meant to say (y+4) not (y-4) at the begining

OpenStudy (freckles):

oh ok

OpenStudy (anonymous):

solution of y=12

OpenStudy (freckles):

\[3^2=\frac{y^2}{y-4} \\ 9=\frac{y^2}{y-4} \\ 9(y-4)=y^2 \\ 9y-36=y^2 \\ 0=y^2-9y+36\]

OpenStudy (freckles):

oops i forgot to replace y-4 with y+4 lol sorry

OpenStudy (freckles):

i was trying to check your quadratic

OpenStudy (freckles):

\[9(y+4)=y^2 \\ 9y+36=y^2 \\ 0=y^2-9y-36\] ok your quadratic looks good

OpenStudy (freckles):

and I guess you factored it maybe (y-12)(y+3)=0

OpenStudy (freckles):

and you knew y=12 would only work because y=-3 would make the insides of the log negative and we can't have that

OpenStudy (freckles):

good job @jamie123